AM3 -Esercitazione 9 A.A. 2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 3 maggio 2004

Esercizio 1

Calcolare

∫∫∫_{D} x^{2} dx dy dz

sull’ellissoide D = \{(x, y, z) ∈ ℝ^{3} : \frac{x^{2}}{a^{2}} + \frac{y^{2}}{b^{2}} + \frac{z^{2}}{c^{2}} ≤ 1\} con a, b, c > 0.

Esercizio 2

Calcolare

∫∫∫_{S} \frac{1}{(y + 1)^{3}} dxdydz

dove

S = {(x, y, z) ∈ ℝ^{3} : 0 < x < 1, 0 < z < 1, 0 < y < x + z}.

Esercizio 3

Dato D = \{(x, y, z) ∈ ℝ^{3} : 0 ≤ x ≤ 1, z + \sqrt{x} ≤ y ≤ 1, 0 ≤ z ≤ 1\} e f : [0, 1] ⟶ ℝ continua, dimostrare che

∫∫∫_{D} f(y) dxdydz = \frac{1}{3} ∫_{0}^{1} t^{3}f(t) dt.